# What drives the growth of black holes: a decade of progress

### v1

@@ 2.1 Scientific context: the broader black-hole research field
1. Facilities: technological advances due to new facilities; e.g., instruments, telescopes, observatories, super computers. |
2. Data: increase in the amount of available data from existing facilities; e.g., larger (or more complete) data samples. |
3. Techniques: new data analysis approaches leading to new insight; e.g., machine learning, Bayesian methods. |
4. Community: group sociology, interactions, sizes, facilitated workshops and discussions, and potential for impact and funding. |
5. Conceptual: new ideas from existing data; e.g., cross disciplinary research, inspiration, focused discussions (e.g., at workshops). |
@@ 3 How does matter accrete onto black holes from kilo-parsec to sub-parsec scales?
To better understand how the gas can be driven inwards to fuel the central SMBH, we describe the main dynamical features of a barred spiral galaxy, the orbital structure, the resonances, and the gravity torques between stars and gas in §3.1. Some fraction of SMBH growth can also be due to the accretion of stars, which is detailed in §3.2. A large part of the progress in the last decade in this research field has been facilitated by more sensitive observations and more extensive data, essentially through increased spatial resolution, brought by new facilities and interferometers. This will be described in §3.3, with ALMA (Atacama Large Millimeter Array) discoveries of molecular tori, kinematically and morphologically decoupled from the main galaxy disk, and with VLTI (Very Large Telescope Interferometer) mid-IR interferometric images, where the main dust emission near the AGN traces out a polar cone geometry. These observations are progressively changing our concepts and paradigm about the dusty torus and the AGN unification model. At even higher spatial resolution (sub-milliarcsec scales probing 0.05 pc), the radio very long baseline interferometry, and the EHT, are now revealing several features of the accretion disk. Our knowledge of the SMBH growth mechanism on these scales is reviewed in §3.4. The small sizes of the AGN emission regions imply that they can vary on day–year time scales. The stronger variations may result from a complete transformation from an optical type 1 AGN (with prominent broad emission lines) to a type 2 AGN (lacking broad emission lines), and these changing-look AGN teach us a lot about the SMBH accretion mechanisms, as described in §3.5. Numerical simulations and new techniques using ever more powerful computers have enlightened these mechanisms in the last decade, as shown in §3.7. Finally, §3.8 summarizes the progress made over the last decade, noting that the future is bright both in the high-spatial resolution and the time domain.
@@ 3.3 Fuelling the AGN: a review of recent observations
The advent of ALMA in 2011 has been instrumental in advancing our empirical understanding of gas inflows down to small ( 10 pc) scales over the last decade (García-Burillo et al., 2014, 2021). It is now possible to trace molecular gas not only in the primary and nuclear bar regions, as expected by the dynamical theory presented in §3.1, but also inside the nuclear spiral structures, which has allowed for the discovery of additional molecular disk components. The latter has been found to be randomly oriented with respect to the rest of the disk, and is kinematically decoupled: we associate this structure to the molecular torus, which is central to the unified AGN model (Urry and Padovani, 1995). The gas fuelling of these molecular tori is thought to be the result of secular evolution, with the nuclear spiral being the last step in the fuelling process (cf §3.1). The molecular tori is within the region where the SMBH dominates the gravitational potential (i.e., at , see Footnote 1 and Fig. 1). The accretion time scale onto the molecular tori within this region is quite short ( 10 Myr), and nuclear bars are more ephemeral than primary bars. However the decoupling of an embedded bar weakens the primary bar, and also the positive torques, outside corotation. This allows the replenishment of the gas disk, so that the molecular torus may be fuelled intermittently.
@@
Conversely, Fig. 9 shows an example of a nuclear spiral, but without the molecular torus in NGC 1097; the latter could have been suppressed by AGN feedback, see §6. Indeed, it has been shown that the cold molecular gas concentration, at scales below 50 pc, is quite high compared to the 200 pc-scale molecular surface density but only below a critical X-ray luminosity (García-Burillo et al., 2024). Above this critical luminosity, the concentration falls by 4 orders of magnitude. This may be attributed to AGN feedback, which depletes the central gas. This result has been obtained for a sample of 64 nearby (D = 7–45 Mpc) disk galaxies including 45 AGN and 19 non-AGN (García-Burillo et al., 2024), where there is sufficient spatial resolution to identify a molecular torus, when present. Additional data is being acquired to improve the source statistics.
@@
The original AGN unification paradigm was based on the existence of a dusty torus whereby the broad line region (BLR) and accretion disk of a Type 1 AGN was obscured if the dusty torus lay along the line of sight towards the observer, such that only the Type 2 AGN features are identified. However, infrared interferometers and in particular VLTI-MATISSE have shown that the dust in the central parts of AGNs are frequently detected in the polar direction, instead of being aligned in the same plane as the putative dusty torus (Asmus et al., 2016; Asmus, 2019; Leftley et al., 2021, 2024). In a revised AGN unification scenario, a conceptual advance driven by new facilities, the obscuration can come both from a dense and geometrically thin dusty disk as well as the base of the polar cone, which is quite clumpy (Isbell et al., 2022, 2023). Both of these structures replace the original geometrically thick torus, where the geometrical thickness corresponded to the relative abundance of Type 1 and Type 2 AGN. This polar dust is most likely ejected within an AGN-driven wind (Alonso-Herrero et al., 2021), starting from the dust sublimation radius (pc size), and forming the border of a hollow cone (Hönig, 2019). The gas is therefore initially predominantly in infall within a geometrically thin disk identified as the molecular torus, but can later be in outflowing in a direction perpendicular to the torus and the accretion disk. Polar cones are ubiquitous in AGN due to the radiation pressure on dust, and their emission is easy to detect because of the high temperature dependence of their flux (Asmus, 2019). Molecular outflows are more difficult to detect along the colder molecular gas, and/or could be dragged by radio jets, which are usually oriented parallel to the wind direction, as sketched in Fig. 10.
@@
Further key progress has been facilitated in this research field by Integral Field Units (IFU) like MUSE on the VLT, which can map the kinematics of the ionized gas across the galaxy (Husemann et al., 2019; García-Bernete et al., 2021). These kinematical studies have been of paramount importance to distinguish between rotational motions of the gas and outflows, that could produce efficient feedback, as discussed in more detail in §6, Outflows accelerated by AGN are commonly observed in the form of coherent, mildly collimated high-velocity gas directed along the AGN ionisation cones, and kinetically powerful jets (García-Bernete et al., 2021). The outflow is comprised of gas in different phases (i.e., it is multi phase), with the highest velocity gas being ionized but where the bulk of the mass in the outflow is in the molecular phase (Fiore et al., 2017). If a jet component is present and lies perpendicular to the disk of the galaxy then it’s impact will be at a maximum, due to a stronger coupling with the ISM gas, even more so if the jet has a low power (i.e., non relativistic) since it can interact with the ISM for a longer duration. This scenario appears to be occurring in NGC 1068, the prototypical nearby Seyfert 2 (García-Burillo et al., 2016, 2019) where the jet-ISM interaction drives a cocoon in the galaxy disk, and produces lateral velocity perturbation perpendicular to the jet direction (Venturi et al., 2021; Girdhar et al., 2022). The gas excitation and broadening of the emission lines is predominantly due to shock ionization. An ionized collimated outflow with symmetric blueshift and redshift may exist in the direction parallel to the jet and wind. Complex situations may occur when the jet and wind are completely aligned with the galaxy disk, entraining ionized and molecular gas outflows, while an oblique ionization cone can also be identified where the ISM is elevated in a fountain by the star formation feedback and illuminated by the central AGN (Husemann et al., 2019). For each outflow “event”, the AGN feedback is not sufficiently energetic or efficient to quench star formation on the galaxy scale, and would have a more local impact. However, when repeated outflows occur, their cumulative effect may well be crucial to slowly quenching star formation; see §6.
@@
High resolution near-IR AGN imaging of NGC 1068 has also been obtained with VLTI-GRAVITY (GRAVITY Collaboration et al., 2020). The interpretation of the results in terms of a circum-nuclear ring has been debated by Gámez Rosas et al. (2022). Finally, combining GRAVITY and MATISSE data has resulted in a coherent multi-wavelength model, where the parsec-scale region is composed of a dust disk plus a polar wind (Leftley et al., 2024). Breakthrough results were also obtained for M 87 using the EHT, not only in imaging the shadow of the SMBH (see §1), but also the radio jet at the base where it is formed and launched, showing that it produces edge-brightened structures connected to the accretion flow and the gravitationally lensed rings (Lu et al., 2023). In Centaurus A, it was also possible to reveal the launching and initial collimation region of the radio jet, down to 10 gravitational radii (Janssen et al., 2021). As in M 87, the radio jet is found to be edge brightened, and its structure is remarkably comparable, when scaled by SMBH mass, leading to the potential identification of a universal profile for SMBHs over a range of scales.
@@ 3.6 Extreme events and quasi periodic eruptions
Extreme quasar variability has been studied as flares superposed on their normal stochastic variability (Graham et al., 2017). These flares occurring in 5 10-5 of the million quasars considered, last typically 3 years. Micro-lensing by stars in foreground galaxies could be one explanation (Lawrence et al., 2016). But the majority could be due to a changing-look event, and explosive stellar-related activity in the accretion disk: super-luminous supernovae, TDEs, and mergers of stellar mass black holes. They can also be due to the interaction of a supermassive black hole with a lower mass companion in an extreme mass-ratio inspiral (EMRI) (Chakraborty et al., 2024), or with a single star (Linial and Metzger, 2023). TDEs can produce indeed spectacular variation, as in the Seyfert 2 galaxy GSN 069, where the X-ray emission increased by more than a factor of 240 in 16 years (Miniutti et al., 2019). After that, the decrease was slow over 8 years, with only a thermal spectrum. However, there were then X-ray quasi-periodic eruptions (QPE), with jumps by two orders of magnitude, over the course of 54 days on time scales of hours. Since GSN 069 is a dwarf galaxy with a 4 10M⊙ SMBH, the class of AGN with more massive SMBHs could have the same eruptions on time scales of months, given the relative relevant sizes. We should note that X-ray bursts are quite common in stellar-mass systems (Galloway et al., 2008), but difficult to identify in SMBHs, although they should also be quite common. A handful of QPE have been detected since 2019 in low-mass AGN-hosts (Arcodia et al., 2024). They are thought to be due to gas accretion from a white dwarf in a very eccentric orbit about the central massive black hole. The orbit of the white dwarf is progressively reduced in radius and excentricity, due to radiation of gravitational waves (King, 2023a). The energy liberated at each passage is then much lower, while the phenomenon becomes a quasi-periodic oscillation, hardy detectable. The constraint of a closely orbiting star with high excentricity favors the low-mass SMBH (King, 2023b).
@@ 4.1.2 X-ray selection
X-ray emission, primarily produced by inverse Compton-scattering of UV photons from the accretion disk by a hot, compact corona (extending to a few , e.g. Fabian et al., 2015), provides an alternative tracer of the AGN accretion power and an efficient means of identifying AGN (see Brandt and Alexander, 2015; Kara and García, 2025, for reviews); see Figs. 1 & 2. A major advantage of X-ray selection of AGN is that other galactic sources of X-rays (such as the population of X-ray binaries throughout the galaxy or hot gas) are relatively weak, resulting in a strong host–to–AGN contrast. Furthermore, higher energy X-rays ( keV) are relatively immune to photoelectric absorption (up to equivalent hydrogen column densities, cm-2) while the absorption of softer X-rays produces a clear spectral signature that enables measurements of the column density. At higher column densities ( cm-2), corresponding to the Compton-thick regime, X-rays of all energies are no longer able to penetrate without being Compton scattered, which substantially suppresses the directly transmitted spectrum, while some photons may be scattered back into the line of the sight with an altered spectrum with a characteristic “Compton hump” peaking at rest-frame energies keV (see Fig. 2). The need for a substantial population of Compton-thick AGN (% of moderate-luminosity AGN) to explain the peak in the overall cosmic X-ray background (CXB) at keV has been well established (e.g. Comastri et al., 1995; Gilli et al., 2007; Ananna et al., 2020). However, an area of significant progress has been the development of models to describe the absorption, line fluorescence, and scattering of X-rays by a realistic obscurer, expanding earlier work adopting toroidal geometries (e.g. Murphy and Yaqoob, 2009; Ikeda et al., 2009; Brightman and Nandra, 2011) to now include clumpy, varying and dynamic distributions (e.g. Tanimoto et al., 2020; Buchner et al., 2019, 2021) that align with our modern picture of accreting SMBH systems (see §3).

### v2

@@ 2.1 Scientific context: the broader black-hole research field
| 1. Facilities: technological advances due to new facilities; e.g., instruments, telescopes, observatories, super computers. |
| 2. Data: increase in the amount of available data from existing facilities; e.g., larger (or more complete) data samples. |
| 3. Techniques: new data analysis approaches leading to new insight; e.g., machine learning, Bayesian methods. |
| 4. Community: group sociology, interactions, sizes, facilitated workshops and discussions, and potential for impact and funding. |
| 5. Conceptual: new ideas from existing data; e.g., cross disciplinary research, inspiration, focused discussions (e.g., at workshops). |
@@ 3 How does matter accrete onto black holes from kilo-parsec to sub-parsec scales?
To better understand how the gas can be driven inwards to fuel the central SMBH, we describe the main dynamical features of a barred spiral galaxy, the orbital structure, the resonances, and the gravity torques between stars and gas in §3.1. Some fraction of SMBH growth can also be due to the accretion of stars, which is detailed in §3.2. A large part of the progress in the last decade in this research field has been facilitated by more sensitive observations and more extensive data, essentially through increased spatial resolution, brought by new facilities and interferometers. This will be described in §3.3, with ALMA (Atacama Large Millimeter Array) discoveries of molecular tori, kinematically and morphologically decoupled from the main galaxy disk, and with VLTI (Very Large Telescope Interferometer) mid-IR interferometric images, where the main dust emission near the AGN traces out a polar cone geometry. These observations are progressively changing our concepts and paradigm about the dusty torus and the AGN unification model. At even higher spatial resolution (sub-milliarcsec scales probing 0.05 pc), the radio very long baseline interferometry, and the EHT, are now revealing several features of the accretion disk. Our knowledge of the SMBH growth mechanism on these scales is reviewed in §3.4. The small sizes of the AGN emission regions imply that they can vary on day–year time scales. The stronger variations may result from a complete transformation from an optical type 1 AGN (with prominent broad emission lines) to a type 2 AGN (lacking broad emission lines), and these changing-look AGN, extreme variability events, and quasar periodic eruptions teach us a lot about the SMBH accretion mechanisms, as described in §3.5 and §3.6. Numerical simulations and new techniques using ever more powerful computers have enlightened these mechanisms in the last decade, as shown in §3.7. Finally, §3.8 summarizes the progress made over the last decade, noting that the future is bright both in the high-spatial resolution and the time domain.
@@ 3.3 Fuelling the AGN: a review of recent observations
The advent of ALMA in 2011 has been instrumental in advancing our empirical understanding of gas inflows down to small ( 10 pc) scales over the last decade (Garc´ıa-Burillo et al., 2014, 2021). It is now possible to trace molecular gas not only in the primary and nuclear bar regions, as expected by the dynamical theory presented in §3.1, but also inside the nuclear spiral structures, which has allowed for the discovery of additional molecular disk components. The latter has been found to be randomly oriented with respect to the rest of the disk, and is kinematically decoupled: we associate this structure to the molecular torus, which is central to the unified AGN model (Urry and Padovani, 1995). The gas fuelling of these molecular tori is thought to be the result of secular evolution, with the nuclear spiral being the last step in the fuelling process (cf §3.1). The molecular tori is within the region where the SMBH dominates the gravitational potential (i.e., at , see Footnote 1 and Fig. 1). The accretion time scale onto the molecular tori within this region is quite short ( 10 Myr), and nuclear bars are more ephemeral than primary bars. However the decoupling of an embedded bar weakens the primary bar, and also the positive torques, outside corotation. This allows the replenishment of the gas disk, so that the molecular torus may be fuelled intermittently.
@@
Conversely, Fig. 9 shows an example of a nuclear spiral, but without the molecular torus in NGC 1097; the latter could have been suppressed by AGN feedback, see §6. Indeed, it has been shown that the cold molecular gas concentration, at scales below 50 pc, is quite high compared to the 200 pc-scale molecular surface density but only below a critical X-ray luminosity (Garc´ıa-Burillo et al., 2024). Above this critical luminosity, the concentration falls by 4 orders of magnitude. This may be attributed to AGN feedback, which depletes the central gas. This result has been obtained for a sample of 64 nearby (D = 7–45 Mpc) disk galaxies including 45 AGN and 19 non-AGN (Garc´ıa-Burillo et al., 2024), where there is sufficient spatial resolution to identify a molecular torus, when present. Additional data is being acquired to improve the source statistics.
@@
The original AGN unification paradigm was based on the existence of a dusty torus whereby the broad line region (BLR) and accretion disk of a Type 1 AGN was obscured if the dusty torus lay along the line of sight towards the observer, such that only the Type 2 AGN features are identified. However, high-resolution infrared observations, including ground-based interferometers (in particular VLTI-MATISSE) have shown that the dust in the central parts of AGNs can often be detected in the polar direction, instead of being aligned in the same plane as the putative dusty torus (Asmus et al., 2016; Asmus, 2019; Leftley et al., 2021, 2024). In a revised AGN unification scenario, a conceptual advance driven by new facilities, the obscuration can come both from a dense and geometrically thin dusty disk as well as the base of the polar cone, which is quite clumpy (Isbell et al., 2022, 2023), with the geometrical thickness from the combination of these obscuring regions corresponding to the relative abundance of Type 1 and Type 2 AGN (Alonso-Herrero et al., 2011; Ramos Almeida et al., 2011; Elitzur, 2012; Ramos Almeida and Ricci, 2017). The polar dust component is most likely ejected (or even formed) within an AGN-driven wind (Alonso-Herrero et al., 2021), starting from the dust sublimation radius (pc size), and forming the border of a hollow cone (Hönig, 2019). The gas is therefore initially predominantly in infall within a geometrically thin disk identified as the molecular torus, but can later be in outflowing in a direction perpendicular to the torus and the accretion disk. However, polar dust is not found in all AGN (for example, observations of two high-Eddington ratio sources show no evidence for polar dust; Drewes et al., 2025) and it is expected to be most commonly identified in sources with accretion properties that are conducive to a long-lived extended outflow (e.g., Venanzi et al., 2020; Garc´ıa-Bernete et al., 2022). Molecular outflows are more difficult to detect along the colder molecular gas, and/or could be dragged by radio jets, which are usually oriented parallel to the wind direction, as sketched in Fig. 10.
@@
Further key progress has been facilitated in this research field by Integral Field Units (IFU) like MUSE on the VLT, which can map the kinematics of the ionized gas across the galaxy (Husemann et al., 2019; Garc´ıa-Bernete et al., 2021; Juneau et al., 2022). These kinematical studies have been of paramount importance to distinguish between rotational motions of the gas and outflows, that could produce efficient feedback, as discussed in more detail in §6, Outflows accelerated by AGN are commonly observed in the form of coherent, mildly collimated high-velocity gas directed along the AGN ionisation cones, and kinetically powerful jets (Garc´ıa-Bernete et al., 2021). The outflow is comprised of gas in different phases (i.e., it is multi phase), with the highest velocity gas being ionized but where the bulk of the mass in the outflow is in the molecular phase (Fiore et al., 2017). If a jet component is present and lies perpendicular to the disk of the galaxy then it’s impact will be at a maximum, due to a stronger coupling with the ISM gas, even more so if the jet has a low power (i.e., non relativistic) since it can interact with the ISM for a longer duration. This scenario appears to be occurring in NGC 1068, the prototypical nearby Seyfert 2 (Garc´ıa-Burillo et al., 2016, 2019) where the jet-ISM interaction drives a cocoon in the galaxy disk, and produces lateral velocity perturbation perpendicular to the jet direction (Venturi et al., 2021; Girdhar et al., 2022). The gas excitation and broadening of the emission lines is predominantly due to shock ionization. An ionized collimated outflow with symmetric blueshift and redshift may exist in the direction parallel to the jet and wind. Complex situations may occur when the jet and wind are completely aligned with the galaxy disk, entraining ionized and molecular gas outflows, while an oblique ionization cone can also be identified where the ISM is elevated in a fountain by the star formation feedback and illuminated by the central AGN (Husemann et al., 2019). For each outflow “event”, the AGN feedback is not sufficiently energetic or efficient to quench star formation on the galaxy scale, and would have a more local impact. However, when repeated outflows occur, their cumulative effect may well be crucial to slowly quenching star formation; see §6.
@@
High resolution near-IR AGN imaging of NGC 1068 has also been obtained with VLTI-GRAVITY (GRAVITY Collaboration et al., 2020). The interpretation of the results in terms of a circum-nuclear ring has been debated by Gámez Rosas et al. (2022). Finally, combining GRAVITY and MATISSE data has resulted in a coherent multi-wavelength model, where the parsec-scale region is composed of a dust disk plus a polar wind (Leftley et al., 2024). Breakthrough results were also obtained for M 87 using the Global Millimeter VLBI Array (GMVA) complemented with ALMA, and Greenland Telescope (GLT) observations to not only image the shadow of the SMBH (see §1), but also the radio jet at the base where it is formed and launched, showing that it produces edge-brightened structures connected to the accretion flow and the gravitationally lensed rings (Lu et al., 2023). In Centaurus A, it was also possible to reveal the launching and initial collimation region of the radio jet, down to 10 gravitational radii (Janssen et al., 2021). As in M 87, the radio jet is found to be edge brightened, and its structure is remarkably comparable, when scaled by SMBH mass, leading to the potential identification of a universal profile for SMBHs over a range of scales.
@@ 3.6 Extreme events and quasi periodic eruptions
Extreme quasar variability has been studied as flares superposed on their normal stochastic variability (Graham et al., 2017). These flares occur in 5 10-5 of the million quasars considered and last typically 3 years. Micro-lensing by stars in foreground galaxies could be one explanation (Lawrence et al., 2016). But the majority could be due to a changing-look event, and explosive stellar-related activity in the accretion disk: super-luminous supernovae, TDEs, and mergers of stellar mass black holes. They can also be due to the interaction of a supermassive black hole with a lower mass companion in an extreme mass-ratio inspiral (EMRI) (Chakraborty et al., 2024), or with a single star (Linial and Metzger, 2023). TDEs can produce indeed spectacular variation, as in the Seyfert 2 galaxy GSN 069, where the X-ray emission increased by more than a factor of 240 in 16 years (Miniutti et al., 2019). After that, the decrease was slow over 8 years, with only a thermal spectrum. However, there were then X-ray quasi-periodic eruptions (QPE), with jumps by two orders of magnitude, over the course of 54 days on time scales of hours. Since GSN 069 is a dwarf galaxy with a 4 10M⊙ SMBH, the class of AGN with more massive SMBHs could have the same eruptions on time scales of months, given the relative relevant sizes. We should note that X-ray bursts are quite common in stellar-mass systems (Galloway et al., 2008), but difficult to identify in SMBHs, although they should also be quite common. A handful of QPE have been detected since 2019 in low-mass AGN-hosts (Arcodia et al., 2024). They are thought to be due to gas accretion from a white dwarf in a very eccentric orbit about the central massive black hole. The orbit of the white dwarf is progressively reduced in radius and excentricity, due to radiation of gravitational waves (King, 2023a). The energy liberated at each passage is then much lower, while the phenomenon becomes a quasi-periodic oscillation, hardy detectable. The constraint of a closely orbiting star with high excentricity favors the low-mass SMBH (King, 2023b).
@@ 4.1.2 X-ray selection
X-ray emission, primarily produced by inverse Compton-scattering of UV photons from the accretion disk by a hot, compact corona (extending to a few , e.g. Fabian et al., 2015), provides an alternative tracer of the AGN accretion power and an efficient means of identifying AGN (see Brandt and Alexander, 2015; Kara and Garc´ıa, 2025, for reviews); see Figs. 1 & 2. A major advantage of X-ray selection of AGN is that other galactic sources of X-rays (such as the population of X-ray binaries throughout the galaxy or hot gas) are relatively weak, resulting in a strong host–to–AGN contrast. Furthermore, higher energy X-rays ( keV) are relatively immune to photoelectric absorption (up to equivalent hydrogen column densities, cm-2) while the absorption of softer X-rays produces a clear spectral signature that enables measurements of the column density. At higher column densities ( cm-2), corresponding to the Compton-thick regime, X-rays of all energies are no longer able to penetrate without being Compton scattered, which substantially suppresses the directly transmitted spectrum, while some photons may be scattered back into the line of the sight with an altered spectrum with a characteristic “Compton hump” peaking at rest-frame energies keV (see Fig. 2). The need for a substantial population of Compton-thick AGN (% of moderate-luminosity AGN) to explain the peak in the overall cosmic X-ray background (CXB) at keV has been well established (e.g. Comastri et al., 1995; Gilli et al., 2007; Ananna et al., 2020). However, an area of significant progress has been the development of models to describe the absorption, line fluorescence, and scattering of X-rays by a realistic obscurer, expanding earlier work adopting toroidal geometries (e.g. Murphy and Yaqoob, 2009; Ikeda et al., 2009; Brightman and Nandra, 2011) to now include clumpy, varying and dynamic distributions (e.g. Tanimoto et al., 2020; Buchner et al., 2019, 2021) that align with our modern picture of accreting SMBH systems (see §3).